# Boosting Mixed-Initiative Co-Creativity in Game Design: A Tutorial

### v1

@@ 4.1.1. Initial Setting
The Initial Setting criterion refers to setting the starting point for the creative process. What will you work on? What problem do you want to solve? What is the initial concept or rule set? Most of the time, this is defined by the human creator when working with systems focused on supporting the user. In these cases, the user already has an established base concept and needs help to better develop it and turn it into an artifact (dos Santos Coutinho and Chaimowicz, 2023; Kreminski et al., 2020). In other cases, as in some approaches that use evolutionary algorithms, the creative process begins with an initial population generated by the system (Gallotta et al., 2023; Liapis et al., 2012a, b). The user can only adapt the content from there, which, of course, substantially limits the creativity of the human.
@@ 4.1.2. Initiative
The Initiative criterion refers to having strong initiative throughout the creative process. As the literature on human creativity and computational creativity shows, there are several ways for the human and the computer to actively participate in the process. This is evidenced by the frameworks of Spoto and Oleynik (Spoto and Oleynik, 2017), Muller et al. (Muller et al., 2020) and Grabe et al. (Grabe et al., 2022), which consider several possible actions for both agents, as mentioned earlier. With this parameter, we intend to analyze who has more initiative between the human and the computer. With so many ways to manifest initiative, it can be difficult to pinpoint where the system falls along this spectrum. We highlight four elements to consider when reflecting on the initiative of each co-creator: quantity of interventions, proactivity, input quality, and variety in the type of actions they can perform. Who is more involved? Can they contribute voluntarily and autonomously? Can their input be considered creative? In what ways can they intervene? We should reflect on these elements together. We might have, for instance, a computational agent that continuously makes suggestions, but these can be considered too random, unoriginal, and not adaptable enough to the partner’s feedback. Furthermore, the system may be limited to making suggestions without a chance to question high-level concepts or make decisions about the artifact being produced. In other cases, the interventions of an agent may be more about assessing what the colleague is doing than contributing creative input.
We emphasize that we do not intend for this framework to be used to perform rigorous classificat ion but rather to serve as a tool for reflection on the mixed-initiative approach one seeks to develop. In human-computer co-creativity, it is ideal that both the human and the computer can actively and regularly intervene during the creative process and can do so in different ways that add value to the generation of the artifact(s). We must then contemplate whether both agents have the same opportunities for action.
@@ 4.1.3. Evaluation
The Evaluation criterion refers to assessing the suggestions and artifact(s) being created. Evaluation can be considered as a subclass of the initiative parameter, but we decided to place it as an independent parameter, as it says a lot about who has more control in the process. Who judges if something is good or not? Who evaluates suggestions and has the power to accept or reject them? One can contribute ideas in the creative process, but do they have space to express their opinion on the ideas of others? If one does not like the output generated in the process, can they criticize it?
When we talk about evaluation, it is not limited to stating whether the artifact is viable and a possible solution to the problem (e.g., assessing if a level is playable and balanced). While this objective assessment is pivotal, it is insufficient in a creative process. It is also necessary to assess whether the artifact is creative, which may involve considerations of how novel, interesting, surprising, and valuable it is (Canaan et al., 2018). This is a more subjective assessment that can potentially lead to disagreements between the participants involved. What happens in the vast majority of current works is that, in the event of disagreements, the human agent is the one who has the power to decide on the course to be taken. In a mixed-initiative co-creative scenario, we expect both the human and the computer to evaluate the generated content, express their likes or dislikes, and try to reach a solution that satisfies both parties.
@@ 4.1.4. Final Decision
The Final Decision criterion refers to deciding when to end the creative process. Typically, the process ends when one or both agents consider that the artifact is finished or the creative problem has been solved. In the current state of the art, it is very unlike that the computational agent has any say on this. The human agent is the only one reflecting if a final solution has been reached and deciding when to conclude the process. The computer can be included in this decision by expressing how it feels about the state of the creative process: “I’m happy with this result. Are you?”. Alternatively, it might disagree with the intention of its co-creator to end the process: “I don’t feel like we have yet reached an ideal solution to our problem.”. Through such dialogue, the process will hopefully end when both partners agree that a final result has been achieved.
This is the last aspect we highlight for researchers to analyze their and other mixed-initiative systems regarding the role that the human and the computer play and the impact they have throughout the collaboration. Reflecting on these four parameters, we get a clearer view of how equal the participants are in the various stages of the process.
@@ 4.1.5. Task Assignment
The Task Assignment criterion ranges from working on separate and distinct tasks to working on the same task. This parameter directly correlates to the XAID (eXplainable AI for Designers) spectrum of domain overlap described by Zhu et al. (Zhu et al., 2018), which distinguishes off-task co-creativity from on-task co-creativity. In on-task co-creative scenarios, agents apply the same tools to the same task. Rezwana and Maher (Rezwana and Maher, 2021, 2023) include task distribution as an interaction component in their model, comprising same task and task divided. The Task Assignment criterion also relates to Kantosalo and Toivonen’s definition of alternating co-creativity in contrast to task-divided co-creativity (Kantosalo and Toivonen, 2016). They emphasize the partners’ concern to satisfy the requirements of both parties in alternating settings. A mixed-initiative tool should be ranked further down the scale of Task Assignment if the human and the computer are working on different tasks so that contributions resulting from one task only slightly affect the other. Higher up the scale must be systems where agents are working on the same task, and the input of one agent directly affects the subsequent intervention of its co-creator. Furthermore, there is another factor to consider besides working on the same task. It is also necessary to reflect on how the participants perform the tasks. They may be working on the same task, but do they perform the same type of actions? Can they make use of the same tools? An illustrative case is the difference between interventions done by designing (where manipulation is more direct) or by setting rules and metrics. The human and computer can both be working on a level design task in terms of layout. However, if one can arrange the level tiles while the other can only handle parameters, they are not working on the same task to the fullest extent.
Admittedly, it may or may not be advantageous to have participants in the creative process working on the same task. In human-human co-creativity, it is also common to have a clear task division, as one person may have more skills for certain aspects of game design and the partner for others. One may be better at drawing and therefore be in charge of game art, while another is concerned with building the characters and storyline of the game. However, if we intend to have a collaborative process in which participants foster the creativity of each other and seek to achieve a result that pleases both, then working on the same task as equals is ideal. It is usual to assume that the computer is incapable of being as creative as the human and to limit the type of tasks it can do because of that. We therefore hope that future approaches may counter this preconception.
@@ 4.1.6. Intervention Pace
The Intervention Pace criterion ranges from long design shifts to immediate feedback. Whether they are working on the same task or not, how long do co-creators wait to give and receive feedback? The exchange of ideas is promoted if the stakeholders can respond almost instantly to additions and changes made to the creative artifact. The turns taken by the participants should be short and adapted to the interventions of each other so that both can express their opinions and preferences throughout the process. In fact, the agents involved may not even have precisely defined shifts, being able to interrupt the partner who is designing if they deem it appropriate. This, of course, implies questions of common sense and mutual respect so that every co-creator has the opportunity to speak without anyone imposing too much on others with their interventions. Rezwana and Maher’s (Rezwana and Maher, 2021, 2023) interaction components of Participation Style and Timing of Initiative relate to the Intervention Pace criterion. Participation Style distinguishes turn-taking from simultaneous contribution of both agents. Timing of Initiative distinguishes planned interventions from spontaneous ones. We consider there is a causal relationship between these two components, as simultaneous collaboration favors spontaneity while turn-taking limits it.
In addition to the timing between interventions, we need to reflect on the size of the content leap they represent. The system may respond immediately to user feedback, but it can be too big of a jump between user-generated and system-generated content. In such a case, the response may skip several design steps ahead without allowing the colleague to intervene halfway through, which limits the discussion and gradual evolution of the artifact throughout the process. The beginning of the creative process should be more focused on defining initial concepts. As such, if a designer is soon confronted with a finished artifact, they feel less room for exploring and discussing ideas, as the creative possibilities have already been narrowed.
In conclusion, a mixed-initiative system with a slow intervention pace is characterized by long design shifts and large content leaps. At the top of the scale are systems with a fast intervention pace, where feedback is immediate and there are no extreme leaps in the content created. Although this parameter should be considered independently of Task Assignment, working on the same task is more likely to allow for a faster intervention pace.
@@ 4.1.7. Explainability
The Explainability criterion ranges from a low level of explainability to a high level of explainability. By explainability, we mean the ability of co-creators to explain their intentions, creative input, and process. Explanatory capabilities are critical for streamlining the collaborative process between humans and computers. As we know, communication is an intrinsic and essential element in human-human collaboration. Despite that, it is often overlooked in human-computer co-creativity. Several authors have identified the need to capacitate computational creativity systems with explainability, identifying requirements and strategies to do so. Zhu et al. (Zhu et al., 2018) propose a subfield of Explainable Artificial Intelligence (XAI) focused on the specific needs of game designers when working with co-creative systems. Another subfield of XAI for contexts of computational creativity is proposed by Llano et al. (Llano et al., 2020) They introduce Explainable Computational Creativity (XCC), specifically concerned with enabling two-way communication between creative systems and human users. They identify four essential design principles for co-creative systems to achieve this: mental models, long-term memory, argumentation, and exposing the creative process. Margarido et al. (Margarido et al., 2022) expand on how these principles can be applied to mixed-initiative co-creativity in game design, suggesting several approaches adapted to distinct use cases. Closely related to explainability, Lin et al. (Lin et al., 2022) identify three non-exhaustive dimensions of communication between humans and computational agents: whether communication is human-initiated or agent-initiated, whether it concerns previous contributions (reflection) or future ones (elaboration), and whether it focuses on a general aspect of the creative task (global scope) or an individual one (local scope).
@@
In order to think about the degree of explainability of our mixed-initiative system, there are several questions we can ask. Are the agents involved in the process able to explain their ideas? Do they have any way of justifying their contributions? If so, in what way(s)? What strategies do they use? What are the means of communication? Is the creative partner able to understand? Can the partner also explain themself and be understood? Is it possible to solve through dialogue the conflicts that may arise? Do both agents have room to further defend ideas that have been rejected? Can they recall past interactions that are relevant to the present discussion? As we answer these questions, it becomes increasingly clear whether the mixed-initiative tool has sufficient explainability. A system with low explainability, and therefore on the bottom of the spectrum, does not have the means for co-creative agents to explain and understand the interventions of each other. A system with high explainability allows for a dialogue very close to what happens in human-human co-creativity. Incidentally, it may even be an improved version of this communication, thanks to audiovisual, memory, and processing capabilities that are unique to computational agents.
@@ 4.2.1. AI Spaceship Generator
Initial setting: In this work, the initial setting is exclusively defined by the computational agent. The creative process begins with the system generating an initial population of spaceships, and the user forcibly works from any of these suggestions or requests the system to generate a new population. The conceptual space predefined by the system is too constrained to allow extensive exploration and cannot be transformed.
Initiative: The computational agent can only intervene by generating several spaceship suggestions, while the user can only select one option and request new generation rounds. On the one hand, the user is the one who decides whether there will be a new round of spaceship generation and from what point (whether from the selected spaceship, a random spaceship, or resetting the population). On the other hand, the user has no way of directly modifying the artifacts or clearly indicating their preferred characteristics, which restricts their creativity. The computational agent only generates more spaceships at the behest of its partner but can create their entire structure, exerting significantly greater control over the visual appearance of the artifacts. Thus, we consider the system to have more initiative.
Evaluation: The human agent evaluates the artifacts by selecting the one they prefer to use as a basis for a new iteration. However, the system can only predict the reasons that lead the human agent to prefer a particular spaceship over others. That is, the evaluation criteria of the human agent are assumed. In theory, the system favors user preferences and adjusts the fitness function accordingly, but this is not guaranteed in practice.
Final decision: The system provides possible solutions, but it is up to the user to decide when they are satisfied with a spaceship, ready to finish the process and import the winning spaceship into the game.
Task assignment: Even though both agents are working on creating spaceship designs, they do not work in the same way. The computational agent intervenes by generating the entire structure of the spaceships, while the human agent can only contribute to the creative process by selecting a spaceship they prefer among those presented to them. This selection will never have a disruptive impact on the subsequent intervention of the computer.
Intervention pace: The agents have clearly defined turns. Even if the user is quick to select one (or none) spaceship to evolve from in the next iteration, there is a long waiting period where the user cannot do anything nor receive feedback until the system completes the new generation. Furthermore, the system responds with new complete spaceship designs, as opposed to short modifications to the selected spaceship or more open-ended spaceship prototypes.
Explainability: There is no explainability. Neither the human nor the computer has the means to express their intentions and justify their contributions.
4.2.2. Sprite Editor for Pixel Art Characters
@@
Initial setting: The human agent is the one who initiates the co-creative process, setting the starting point to generate other poses for the character. The computer cannot generate meaningful content without sufficiently comprehensive input from the human partner. Furthermore, if the initial design of the human agent is scarcely recognizable to the trained model, the quality of the contributions of the computational agent is compromised.
Initiative: The human agent has the stronger initiative. They can directly edit any pose view, including those in the suggestion slots. By being able to draw and promptly modify any artifact, the human agent is granted more creative freedom. On the other hand, the computer can only intervene upon user request. As mentioned by the authors, it lacks proactivity to make independent contributions. The computational agent provides suggestions but cannot manipulate the artifacts directly. Moreover, it always tries to match the design provided by the human partner, avoiding the suggestion of images that are conceptually different from the source. Thus, despite generating new pose views, the input from the computational agent lacks creativity.
Evaluation: The system provides suggestions that align with the input from the user without having an opinion about what is being drawn. The evaluation process is entirely carried out by the human agent, who determines whether to accept or ignore suggestions and modify the content if they deem it necessary for improvement.
Final decision: The human designer is the one who decides if and when they are satisfied with the character sprites.
Task assignment: Both agents are engaged in the creation of sprites for a pixel art character. The human agent has the ability to draw the character from scratch, directly manipulate any pose view, and select each image that will serve as input for a new suggestion. On the other hand, the computational agent can generate sprites for the character but can only present them as suggestions. Nevertheless, we can assert that both agents are able to generate content within the same domain.
Intervention pace: The participants involved in the co-creative process can engage in a fast-paced interaction loop. Despite having well-defined turns for each participant (with the system only intervening upon user request), the response is immediate. In most cases, the suggestions from the system do not represent an extensive leap in content and are closely aligned with the user-generated content. Since the human agent has more control in the process, they can generate from none to a substantial amount of new content between the interventions of the computational agent. Although the turns are short, the system cannot intervene spontaneously.
Explainability: There is no explainability.
4.2.3. Sentient Sketchbook
@@
Initial setting: The conceptual space is already predefined by the system. The tool is designed to generate maps for strategy games, meaning that the available tiles for constructing the map and the fitness dimensions used are already established for this purpose. Consequently, the human agent is constrained to work within a specific game style and towards a predetermined objective. As for initiating the co-creative process, the computational agent immediately starts offering suggestions, even if the human agent has not yet begun editing the map. On the other hand, the human designer can start drawing right away, triggering an instant change in the system’s suggestions and thereby establishing the starting point for shaping the map. Taking all these factors into account, we consider that the initial setting criterion leans slightly towards the computer end of the spectrum due to the conceptual space being fully predefined.
Initiative: The human agent contributes by directly editing the map sketch and selecting suggestions. The computational agent intervenes by consistently suggesting alternative map designs but cannot directly edit the main sketch. Both agents are highly proactive and engage multiple times throughout the creative process. The human partner can continuously add tiles and analyze suggestions, while the computer willingly responds with new creative input after each contribution made by its counterpart. Thus, both agents are deeply involved in the process and attentive to each other’s feedback. In this criterion, the user slightly stands out due to the ability to directly manipulate the map sketch.
Evaluation: Both evaluate the content being created. The human agent self-assesses the content they are creating and identifies areas for improvement over time. Simultaneously, they evaluate the map suggestions of the computational agent, considering whether any are superior to the current sketch and possibly choosing to replace it or draw inspiration from certain features they appreciate in the suggestions. On the other hand, the computer tests the playability of the main map sketch and evaluates it based on gameplay properties related to game pace and player balance. It does the same to generate map designs and choose the best ones to present to the user. Additionally, half of the map designs generated by the system are evaluated in terms of novelty, and the system selects the most distinct maps to include in the suggestions. There is, however, a significant difference in the evaluation process carried out by the two agents: the system can only disclose its opinion to the human partner regarding gameplay characteristics. It cannot express whether it finds a particular map creative or not. Furthermore, regardless of what the computational agent informs the user about its evaluation of each of the map properties, only the human designer has the authority to accept or reject suggestions. Therefore, the human has more evaluation power than the computer.
Final decision: The human agent determines when they believe the map sketch is finalized and ready to be exported. Subsequently, the computational agent generates a detailed view of the map. The human partner cannot directly edit the detailed version of the map but can choose their preferred option from the viewing modes provided by the computer or request the generation of a new view. Thus, the human has the final say over the structure of the level but has less control over its final look.
Task assignment: The agents are collaborating on the shared task of creating map designs for a strategy game. The input from the human agent directly affects each subsequent intervention of the computational agent, whereas the contributions of the computational agent may or may not impact the next intervention by the human agent. Although the computer cannot edit the main sketch and select suggestions, both agents can willingly construct map designs.
Intervention pace: The co-creators engage in a fast intervention pace, with short turns involving the addition of small portions of content over time. In this case, the human agent edits a single tile (or a small area) at a time, and the computational agent intervenes with each change made. The content leap between the main map sketch and the computer’s suggestions is also not too substantial. There is only a larger gap observed at the beginning of the co-creative process as the system generates suggestions with sufficient content to be playable. In essence, the agents provide immediate and spontaneous feedback.
Explainability: The explainability is severely limited. The computational agent makes a feeble attempt to justify why its suggestions may be preferable by displaying their respective fitness scores compared to the main map sketch. Thus, in the case of suggestions aimed at optimizing fitness scores, the human agent has a means to understand some of the improvements the suggestions offer to gameplay. However, these properties solely describe functional characteristics and provide no insight into the creativity of the map design. Conversely, the human designer lacks any means to explain their intentions and preferences.
4.2.4. RL Brush
@@
Initial setting: The conceptual space is predefined by the system. Although the RL Brush tool can be applied in theory to any type of tile-based game, this version of the tool was specifically built to design levels for the game Sokoban. As a result, only the tiles necessary for this game are made available. The human agent is restricted to working with these tiles and cannot add new ones or change the conceptual space. The system also takes the first step in initiating the creative process. It invariably presents the user with an empty map containing a player tile at the center, along with four initial suggestions. Even though the computer starts the co-creative process by only inserting a tile (rather than presenting a complete level), it can still slightly influence the direction in which the level will develop. However, since the human agent can immediately start editing the level, they can easily change this initial configuration, triggering new suggestions from the computational agent and taking control of the process.
Initiative: Similar to the Sentient Sketchbook, the human agent directly impacts the creative artifact throughout the process by being able to manually edit the level grid or choose a suggestion to replace it. The computational agent provides suggestions whenever the human partner makes a modification, but cannot directly edit the current level. This approach has the additional characteristic that the user can control the performance of the system, granting them the power to limit the quality and scope of its suggestions. Consequently, both agents proactively contribute creative input, but the human designer retains more control over the process.
Evaluation: The evaluation is mostly done by the human agent. They not only self-assess the content they create but also evaluate the suggestions provided by the computational agent, deciding whether any of them are good enough to be accepted. As for the computational agent, the internal evaluation it performs to generate suggestions remains unclear. We lack insight into the considerations that the AI models of the system have regarding the modifications they propose and how these suggestions may enhance the current level. Even if the system does conduct some form of evaluation, it cannot express it. Therefore, it is the human agent who guides the level design process, retaining complete control over the desired outcome for the creative artifact.
Final decision: It is up to the human agent to decide when they are content with the current level and ready to finish the creative process.
Task assignment: Both agents are working on the shared task of designing a level for the Sokoban game. The contributions made by the human agent directly influence the suggestions presented by the computational agent. Conversely, the creative input from the computational agent may or may not affect the creative process carried out by the human co-creator. Lastly, even though the system cannot directly edit the grid of the current level, both agents contribute by adding or altering tiles.
Intervention pace: The intervention pace of the computational agent is highly dependent on the values of the parameters “step” and “tool radius”. When the “step” value is low and the “tool radius” is small, the feedback of the system is nearly instantaneous. However, if these values are increased, the system’s response time also increases. Nevertheless, we consider this waiting time to be reasonable. Both the human and the computer edit one or a few tiles at a time, resulting in small content leaps. Consequently, the agents are not overwhelmed by a large amount of new information all at once, allowing them to gradually build the level and merge their ideas more easily.
Explainability: There is no explainability.
4.2.5. miWFC

### v2

@@ 4.1.1. Initial Setting
Definition:
Setting the starting point for the creative process.
Key points to consider:
What will you work on? What problem do you want to solve? What is the initial concept or rule set? Think about who can make these decisions. The human? The computer? Both? Can they do it to the same extent?
Illustration:
Often, the initial setting is defined by the human creator when working with systems focused on supporting the user. In these cases, the user already has an established base concept and needs help to better develop it and turn it into an artifact (Coutinho and Chaimowicz, 2024; Kreminski et al., 2020). In other cases, as in some approaches that use evolutionary algorithms, the creative process begins with an initial population generated by the system (Gallotta et al., 2023; Liapis et al., 2012a, b). The user can only adapt the content from there, which, of course, substantially limits the creativity of the human.
Additional remarks:
@@ 4.1.2. Initiative
Definition:
Having strong initiative throughout the creative process.
Key points to consider:
We highlight four elements to consider when reflecting on the initiative of each co-creator: quantity of interventions, proactivity, input quality, and variety in the type of actions they can perform. Who is more involved? Can they contribute voluntarily and autonomously? Can their input be considered creative? In what ways can they intervene? We should reflect on these elements together.
Illustration:
We might have, for instance, a computational agent that continuously makes suggestions, but these can be considered too random, unoriginal, and not adaptable enough to the partner’s feedback. Furthermore, the system may be limited to making suggestions without a chance to question high-level concepts or make decisions about the artifact being produced. In other cases, the interventions of an agent may be more about assessing what the colleague is doing than contributing creative input.
Additional remarks:
As the literature on human creativity and computational creativity shows, there are several ways for the human and the computer to actively participate in the process. This is evidenced by the frameworks of Spoto and Oleynik (Spoto and Oleynik, 2017), Muller et al. (Muller et al., 2020) and Grabe et al. (Grabe et al., 2022), which consider several possible actions for both agents. With so many ways to manifest initiative, it can be difficult to pinpoint where the system falls along this spectrum. In human-computer co-creativity, it is ideal that both the human and the computer can actively and regularly intervene during the creative process and can do so in different ways that add value to the generation of the artifact(s). We must then contemplate whether both agents have the same opportunities for action.
@@ 4.1.3. Evaluation
Definition:
Assessing the suggestions and artifact(s) being created.
Key points to consider:
Who judges if something is good or not? Who evaluates suggestions and has the power to accept or reject them? One can contribute ideas in the creative process, but do they have space to express their opinion on the ideas of others? If one does not like the output generated in the process, can they criticize it?
Additional remarks:
When we talk about evaluation, it is not limited to stating whether the artifact is viable and a possible solution to the problem (e.g., assessing if a level is playable and balanced). While this objective assessment is pivotal, it is insufficient in a creative process. It is also necessary to assess whether the artifact is creative, which may involve considerations of how novel, interesting, surprising, and valuable it is (Boden, 2004; Canaan et al., 2018). This is a more subjective assessment that can potentially lead to disagreements between the participants involved. What happens in the vast majority of current works is that, in the event of disagreements, the human agent is the one who has the power to decide on the course to be taken. In a mixed-initiative co-creative scenario, we expect both the human and the computer to evaluate the generated content, express their likes or dislikes, and try to reach a solution that satisfies both parties.
@@ 4.1.4. Final Decision
Definition:
Deciding when to end the creative process.
Key points to consider:
Typically, the process ends when one or both agents consider that the artifact is finished or the creative problem has been solved. Identify who can make this decision or who has more authority over it.
Additional remarks:
In the current state of the art, it is very unlikely that the computational agent has any say on this. The human agent is the only one reflecting if a final solution has been reached and deciding when to conclude the process. The computer can be included in this decision by expressing how it feels about the state of the creative process: “I’m happy with this result. Are you?”. Alternatively, it might disagree with the intention of its co-creator to end the process: “I don’t feel like we have yet reached an ideal solution to our problem.”. Through such dialogue, the process will hopefully end when both partners agree that a final result has been achieved.
@@ 4.1.5. Task Assignment
Definition:
Ranging from working on separate and distinct tasks to working on the same task.
Key points to consider:
A mixed-initiative tool should be ranked further down the scale of Task Assignment if the human and the computer are working on different tasks, so that contributions resulting from one task only slightly affect the other. Higher up the scale must be systems where agents are working on the same task, and the input of one agent directly affects the subsequent intervention of its co-creator. It is also necessary to reflect on how the participants perform the tasks. They may be working on the same task, but do they perform the same type of actions? Can they make use of the same tools?
Illustration:
An illustrative case is the difference between interventions done by designing (where manipulation is more direct) or by setting rules and metrics. The human and the computer can both be working on a level design task in terms of layout. However, if one can arrange the level tiles while the other can only handle parameters, they are not working on the same task to the fullest extent.
Additional remarks:
This parameter directly correlates to the XAID (eXplainable AI for Designers) spectrum of domain overlap described by Zhu et al. (Zhu et al., 2018), which distinguishes off-task co-creativity from on-task co-creativity. In on-task co-creative scenarios, agents apply the same tools to the same task. Rezwana and Maher (Rezwana and Maher, 2021, 2023) include task distribution as an interaction component in their model, comprising same task and task divided. The Task Assignment criterion also relates to Kantosalo and Toivonen’s definition of alternating co-creativity in contrast to task-divided co-creativity (Kantosalo and Toivonen, 2016). They emphasize the partners’ concern to satisfy the requirements of both parties in alternating settings.
Admittedly, it may or may not be advantageous to have participants in the creative process working on the same task. In human-human co-creativity, it is also common to have a clear task division, as one person may have more skills for certain aspects of game design and the partner for others. One may be better at drawing and therefore be in charge of game art, while another is concerned with building the characters and storyline of the game. However, if we intend to have a collaborative process in which participants foster the creativity of each other and seek to achieve a result that pleases both, then working on the same task as equals is ideal. It is usual to assume that the computer is incapable of being as creative as humans and to limit the type of tasks it can do because of that. We therefore hope that future approaches may counter this preconception.
@@ 4.1.6. Intervention Pace
Definition:
Ranging from long design shifts to immediate feedback.
Key points to consider:
Whether they are working on the same task or not, how long do co-creators wait to give and receive feedback? How much of a leap in content is there between interventions? A mixed-initiative system with a slow intervention pace is characterized by long design shifts and large content leaps. At the top of the scale are systems with a fast intervention pace, where feedback is immediate and there are no extreme leaps in the content created.
Illustration:
Short shifts do not necessarily mean short content leaps. A system may respond immediately to user feedback, but it can be too big of a jump between user-generated and system-generated content. In such a case, the response may skip several design steps ahead without allowing the colleague to intervene halfway through, which limits the discussion and gradual evolution of the artifact throughout the process.
Additional remarks:
Although this parameter should be considered independently of Task Assignment, working on the same task is more likely to allow for a faster intervention pace. The exchange of ideas is promoted if the stakeholders can respond almost instantly to additions and changes made to the creative artifact. The turns taken by the participants should be short and adapted to the interventions of each other so that both can express their opinions and preferences throughout the process. In fact, the agents involved may not even adhere to precisely defined shifts, being able to interrupt the partner who is designing if they deem it appropriate. This naturally raises considerations of practical judgment and sensibility towards one another, so that every co-creator has the opportunity to speak without anyone imposing too much on others with their interventions. In addition, there shouldn’t be large gaps between content. If a designer is soon confronted with a finished artifact, they feel less room for exploring and discussing ideas, as the creative possibilities have already been narrowed.
Rezwana and Maher’s (Rezwana and Maher, 2021, 2023) interaction components of Participation Style and Timing of Initiative relate to this criterion. Participation Style distinguishes turn-taking from simultaneous contribution of both agents. Timing of Initiative distinguishes planned interventions from spontaneous ones. We consider there is a causal relationship between these two components, as simultaneous collaboration favors spontaneity while turn-taking limits it.
@@ 4.1.7. Explainability
Definition:
Ranging from a low level of explainability to a high level of explainability. By explainability, we mean the ability of co-creators to explain their intentions, creative input, and process.
Key points to consider:
@@
Are the agents involved in the process able to explain their ideas? Do they have any way of justifying their contributions? If so, in what way(s)? What strategies do they use? What are the means of communication? Is the creative partner able to understand? Can the partner also explain themself and be understood? Is it possible to solve through dialogue the conflicts that may arise? Do both agents have room to further defend ideas that have been rejected? Can they recall past interactions that are relevant to the present discussion? As we answer these questions, it becomes increasingly clear whether the mixed-initiative tool has sufficient explainability. A system with low explainability, and therefore at the bottom of the spectrum, does not have the means for co-creative agents to explain and understand the interventions of each other. A system with high explainability allows for a dialogue very close to what happens in human-human co-creativity. Incidentally, it may even be an improved version of this communication, thanks to audiovisual, memory, and processing capabilities that are unique to computational agents.
Additional remarks:
Explanatory capabilities are critical for streamlining the collaborative process between humans and computers. As we know, communication is an intrinsic and essential element in human-human collaboration. Despite that, it is often overlooked in human-computer co-creativity. Several authors have identified the need to capacitate computational creativity systems with explainability, identifying requirements and strategies to do so. Zhu et al. (Zhu et al., 2018) propose a subfield of Explainable Artificial Intelligence (XAI) focused on the specific needs of game designers when working with co-creative systems. Another subfield of XAI for contexts of computational creativity is proposed by Llano et al. (Llano et al., 2020). They introduce Explainable Computational Creativity (XCC), specifically concerned with enabling two-way communication between creative systems and human users. They identify four essential design principles for co-creative systems to achieve this: mental models, long-term memory, argumentation, and exposing the creative process. Margarido et al. (Margarido et al., 2022) expand on how these principles can be applied to mixed-initiative co-creativity in game design, suggesting several approaches adapted to distinct use cases. Closely related to explainability, Lin et al. (Lin et al., 2022) identify three non-exhaustive dimensions of communication between humans and computational agents: whether communication is human-initiated or agent-initiated, whether it concerns previous contributions (reflection) or future ones (elaboration), and whether it focuses on a general aspect of the creative task (global scope) or an individual one (local scope).
@@ 4.2.1. Initial Setting
AI Spaceship Generator (SG): In this work, the initial setting is exclusively defined by the computational agent. The creative process begins with the system generating an initial population of spaceships, and the user forcibly works from any of these suggestions or requests the system to generate a new population. The conceptual space predefined by the system is too constrained to allow extensive exploration and cannot be transformed.
@@
Sprite Editor for Pixel Art Characters (SE): The human agent is the one who initiates the co-creative process, setting the starting point to generate other poses for the character. The computer cannot generate meaningful content without sufficiently comprehensive input from the human partner. Furthermore, if the initial design of the human agent is scarcely recognizable to the trained model, the quality of the computational agent’s contributions is compromised.
@@
Sentient Sketchbook (SS): The conceptual space is already predefined by the system. The tool is designed to generate maps for strategy games, meaning that the available tiles for constructing the map and the fitness dimensions used are already established for this purpose. Consequently, the human agent is constrained to work within a specific game style and towards a predetermined objective. As for initiating the co-creative process, the computational agent immediately starts offering suggestions, even if the user has not yet begun editing the map. On the other hand, the user can start drawing right away, triggering an instant change in the system’s suggestions and thereby establishing the starting point for shaping the map. Taking all these factors into account, we consider that the initial setting criterion leans slightly towards the computer end of the spectrum due to the conceptual space being fully predefined.
@@
RL Brush (RB): The conceptual space is predefined by the system. Since this version of the tool was specifically built to design levels for the Sokoban game, only the tiles necessary for this game are made available. The human agent is restricted to working with these tiles and cannot add new ones or change the conceptual space. The system also takes the first step in initiating the creative process. It invariably presents the user with an empty map containing a player tile at the center, along with four initial suggestions, which can slightly influence the direction in which the level will develop. However, since the human agent can immediately start editing the level, they can easily change this initial configuration, triggering new suggestions from the computational agent and taking control of the process.